OpenStudy (anonymous):

Can I get some help answering 5-3a-4/5=7-2a/2

OpenStudy (blurbendy):

5 - 3a -4/5 = 7 -2a/2 -3a +21/5 = 7 - 2a/2 -3a = -21/5 + 7 -2a/2 -3a = 14/5 - 2a/2 -3a + 2a/2 = 14/5 -2a = 14/5 a = -7/5

OpenStudy (anonymous):

answer is suppose to be -23/4

OpenStudy (blurbendy):

no

OpenStudy (blurbendy):

substitute -7/5 in for a into the original equation, and you will get a true statement

OpenStudy (blurbendy):

if you do that for -23/4 you get a false statement.